Liquid octane CH

3

CH

2

6

CH

3

 will react with gaseous oxygen O

2

 to produce gaseous carbon dioxide CO

2

 and gaseous water H

2

O

. Suppose 1.14 g of octane is mixed with 7.1 g of oxygen. Calculate the maximum mass of carbon dioxide that could be produced by the chemical reaction. Round your answer to 3

significant digits. 



1
Expert's answer
2021-11-16T14:03:06-0500

Solution:

The molar mass of octane is 114 g/mol

The molar mass of O2 is 32 g/mol

 

Determine moles of each reactant:

(1.14 g octane) × (1 mol octane / 114 g octane) = 0.010 mol octane

(7.1 g O2) × (1 mol O/ 32.00 g O2) = 0.222 mol O2


Balanced chemical equation:

2CH3(CH2)6CH3(l) + 25O2(g) → 16CO2(g) + 18H2O(g)


According to the chemical equation above:

2 mol of octane reacts with 25 mol of O2

Thus, 0.010 mol of octane reacts with:

(0.010 mol octane) × (25 mol O2 / 2 mol octane) = 0.125 mol O2

However, initially there is 0.222 mol of O2 (according to the task).

Therefore, octane acts as limiting reagent and O2 is excess reagent.

 

According to stoichiometry:

2 mol of octane produces 16 mol of CO2

Thus, 0.010 mol of octane produces:

(0.010 mol octane) × (16 mol CO2 / 2 mol octane) = 0.08 mol CO2


The molar mass of CO2 is 44 g/mol

Therefore,

(0.08 mol CO2) × (44 g CO2 / 1 mol CO2) = 3.52 g CO2


Answer: The mass of carbon dioxide (CO2) is 3.52 grams